OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.

OpenStudy (anonymous):

@zepdrix can you please help me.

zepdrix (zepdrix):

The Slope Intercept Form of a line is given by:\[\Large y=mx+b\]Where \(\Large m\) is the slope of our line. And \(\Large b\) is our y-intercept. So in our example it looks like we have \(\large m=2\) and \(\large b=-3\) right?

OpenStudy (anonymous):

In slope intercept form eq of line is y=mx+c where c= y intercept=3 m= slope=2 So y=mx+c becomes y=2x+3